Kingbird, Inc. manufactures souvenirs and sells the products to souvenir shops across Canada. Lisa McKay is the new owner, and is concerned about the low margins. She would like to find a way to improve the company’s profitability. The accountant provides her the following financial information: sales are $306,000, of which 60% is cost of goods sold. Cost of goods sold consists of direct materials (20%), direct labour (30%), and fixed manufacturing overhead (50%). Operating expenses consist of variable expenses (40%) and fixed expenses (60%). Kingbird pays a 40% tax rate and the net income is $14,100. To reduce the company’s operating risks, McKay would like to review the company’s operations from another perspective. She would like to know how much the company needs to generate in order to break-even. Based on the current cost structure, how sensitive is the profit to a sale volume increase of 5%.

Prepare an income statement in the gross margin format. Kingbird, Inc. Gross Margin Income StatementPrepare an income statement in the contribution margin format. Kingbird, Inc. Contribution Margin Income StatementWhat is the break even in sales? Round contribution margin percentage to 4 decimal places e g. O 4687 and fi al answer to decima Places e 125 Sales break-even$ LINK TO TEXT LINK TO TEXT LINK TO TEXT LINK TO TEXT At the current sales level, what is the margin of safety and the degree of operating leverage? (Round degree of operating leverage to 2 decimal places, e.g. 1.25.) Margin of Safety Degree of Operating Leverage LINK TO TEXT LINK TO TEXT LINK TO TEXT LINK TO TEXT
Answer #1
Cost of goods sold and its division:
Total sales 306000
Cost of goods sold as % of sales 60%
Cost of goods sold in $ 183600
Material (20% of 183600) 36720
Labour (183600*30%) 55080
Fixed manufacturing Oh (183600*50%) 91800
Computation of Operating expense and its allocation;
Net income 14100
Tax rate 40%
Net income before tax (14100/60%) 23500
Gross Profit (sales -Cost of goods sold) 122400
(306000-183600)
Operating expense (Gross profit- Net income before tax) 98900
Variable operating expenses (98900*40%) 39560
Fixed operating expense (98900*60%) 59340
Gross margin income Statement:
Sales revenue 306000
Less: Cost of goods sold
Material 36720
Labour 55080
Manufacturing Oh 91800
Total cost of goods ssold 183600
Gross margin     122400
Less: Operating expense
Variable expenses 39560
Fixed expenses 59340
Total Operating expenses 98900
Net income before tax 23500
Less: Tax @ 40% 9400
After tax net income 14100
Contribution Margin Income Statement:
Sales revennue 306000
Less: variable cost
Material 36720
labour 55080
variable operating expense 39560
Total Variable cost 131360
Contribution margin    174640
Less: Fixed cost
Fixed manufacturing oH 91800
Fixed operating expenses 59340
Total fixed cost 151140
Net income before tax 23500
Less: Tax @ 40% 9400
After tax income 14100
CM ratio = Contribution / Sales *100
174640 /306000 *100 = 57.07%
Break even sales = Total fixed cost / CM ratio
151140 /57.07 % = $ 264833
margin of safety = Actual sales - Break even sales
306000 -264833 = 41167
Degree of Operating leverage = Contribution/ Net income before tax
174640 /23500 = 7.43 times